OpenStudy (anonymous):

sin(x+10deg)=cos3x then x=?

OpenStudy (anonymous):

can you help me ?

hartnn (hartnn):

u know the identity sin(90-A)=cos A ?

hartnn (hartnn):

so write cos 3x as sin (90-3x) now u have sin(x+10) = sin (90-3x) so x+10 = 90-3x can u solve this for x ?

OpenStudy (anonymous):

Note that, \[x + 10 + 2n \pi = 90 - 3x\]

OpenStudy (anonymous):

i get 20 :)) thank you so much

hartnn (hartnn):

yup, its 20, do u need general solution in terms of n?

OpenStudy (anonymous):

yes yes :)

hartnn (hartnn):

x+10+2nπ=90−3x 4x=80-2nπ x=20-nπ/2

OpenStudy (anonymous):

oh i get it. thanks but can i have 1 more question?

OpenStudy (anonymous):

If cosA= -15/17 and A is in quadrant III, find cos 1/2A